OpenStudy (anonymous):

Could someone check my solution to this? A shell explodes on the surface of horizontal ground. Earth is scattered in all directions with varying velocities. Show that particles of earth with initial speed v landing a distance r from the centre of explosion will do so at times t given by \[\frac{1}{2}g^2t^2=v^2 \pm \sqrt{v^4-r^2g^2}\]

OpenStudy (anonymous):

http://gyazo.com/50f04a7f09df3fa4192bceae0367fb74

OpenStudy (anonymous):

I half-reached the solution, but the plus-minus killed my method, as in general \[\sin^2x \ne \cos^2x\] Where did I make the logical error?

OpenStudy (zehanz):

But will the particles not also have all kinds of vertical angles \(\theta\) at which they are being scattered? Then \(v \sin \theta\) would determine the time needed to get to the ground again.

OpenStudy (anonymous):

Of course, which is why I have left the angle as arbitrary. However, that makes little inroads into the question.

OpenStudy (zehanz):

I've got to the part that says \(v^4-g^2......\). Where did that come from?

OpenStudy (anonymous):

That's \[v^4-g^2r^2\[, the thing inside the square root

OpenStudy (anonymous):

\[v^4-g^2r^2\]

OpenStudy (anonymous):

Substituting in what I found for r.

OpenStudy (zehanz):

But inside the square root in your question it was \(v^2\)...

OpenStudy (anonymous):

That was a typo- sorry!!

OpenStudy (anonymous):

http://gyazo.com/c7996dccd0b5b0a9d95d7161217a4872 Here's the original question as I'm evidently incompetent to duplicate it correctly.

OpenStudy (zehanz):

OK, I'll go on with it...

OpenStudy (anonymous):

Thanks

OpenStudy (zehanz):

I get everything the same as you:\[\sqrt{v^4-g^2r^2}=\sqrt{v^4\cos^22w}=v^2|\cos2w|=\pm v^2\cos2w\]So\[v^2 \pm v^2\cos2w=v^2(1 \pm \cos2w)=v^2(1 \pm(1-2\sin^2w))\]The plus or minus leads to \(2v^2\cos^2w\) and \(2v^2\sin^2w\). I do not see any errors in your calculation. Could it have something to do with the value of w? If w > 45 deg, then r becomes smaller again, and cos2w (in the first line of this reply) becomes negative.

OpenStudy (anonymous):

The problem is that this is supposed to be equal to \[0.5g^2t^2\] Which is \[2v^2\sin^2(w)\] However, one of those solutions was\[2v^2\cos^2(w)\], which is obviously not the same.

OpenStudy (zehanz):

Yeah, that was my problem too...